# Kishu: Time-Traveling for Computational Notebooks

### v1

@@ 1. Introduction
Oftentimes, during a workflow, users would like to revert changes made to the session state (i.e., time-travel), such as to undo a modification (e.g., restore a dropped column of a dataframe (StackOverflow, 2024)), restoring an overwritten variable (Groups, 2024), or perform reverse debugging (Brachmann and Spoth, 2020). Unfortunately, unlike program debuggers (e.g., gdb) (Phang et al., 2013; Barr and Marron, 2014; GDB, 2024), relational databases (e.g., PITR in PostgreSQL and MySQL (Group, 1996a; Oracle, 2024b)) or interactive data systems (Dunne et al., 2012; Crotty et al., 2015; Kraska, 2021) which support time-traveling to past program states, existing notebook systems do not natively keep track of past session states: cell executions cannot be undone, e.g., the user cannot ’un-drop’ a dataframe column. If the user executes a cell that alters the session state, a common approach to restore the previous state would be to restart the kernel and then (painstakingly) re-run past cells in the correct order. While code versions can be saved using tools such as Git (Git, 2024b) or native commands (e.g., Jupyter’s %checkpoint (Team, 2023b) [^1] Despite its name, %checkpoint only stores cell code and not objects in the state.) to simplify identifying cells to rerun for restoration, cell reruns can still be time-consuming (e.g., re-training an ML model) and/or result in incorrect restoration (e.g., random train-test splits). Another approach is for the user to periodically checkpoint the session state (e.g., memory dump (CRIU, 2023; Ansel et al., 2009) or session state serialization (Foundation, 2023h)) to storage or a managed database (e.g., KV-store (Team, 2023c)). Then, users can load an appropriate checkpoint file to restore the session state. However, performing session checkpointing and restoration using these tools is limiting: checkpointing can incur prohibitive costs (§ 7.3, § 7.4) and may fail on certain workloads (e.g., GPU (cri, [n.d.])), and restoration can either (1) only be (inefficiently) performed from scratch, requiring completely loading a checkpoint file (Foundation, 2023h) and/or killing the current kernel (CRIU, 2023), or (2) may be incorrect, breaking inter-variable relations (Team, 2023c).
@@
We propose Kishu, a notebook system that enables time-traveling between session states: as the user executes cells, Kishu tracks the session state evolution while writing per-cell incremental checkpoints containing differing data between successive states (i.e., the state delta) for returning to any past state via an incremental checkout later. Kishu pursues three challenging goals—Delta-Efficient Checkpoint: Kishu aims to minimize incremental checkpointing overhead by exploiting the small per-cell deltas typical of data science workflows (§ 7.6.1), but also avoid high detection overhead in the face of complex access patterns and inter-variable dependencies. Correct & Non-intrusive Checkout: Kishu aims to restore past states in the same session non-intrusively by leveraging existing objects in the kernel (that don’t need updating) to minimize data loading costs, while still guaranteeing checkout accuracy as if it completely loaded a checkpoint file. Generalizability: Kishu aims to support checkpointing/checkout for almost all notebook libraries (and/or use cases), of which there is a large variety, e.g., notebooks can perform distributed computing (e.g., Spark (Zaharia et al., 2010)) or move data off-CPU (e.g., GPUs (Foundation, 2024a)). If Kishu can achieve these goals, Kishu will allow users to undo almost any executed cell that undesirably modifies the state as if it never occurred by quickly checking out to the pre-execution state at the cost of minimal workflow overhead.
@@
Our core idea for achieving the aforementioned goals is to capture the session state delta with low overhead, but at a sufficiently high granularity using information exclusively available at the application level; this will allow us to perform generalizable, correct, and efficient time-traveling, as follows:
@@
Second, for correct incremental checkout, Kishu identifies the difference between the current and target session state at the aforementioned Co-variable granularity via state divergence according to the checkpoint graph. Then, it replaces (only) Co-variables that need updating in the session state by loading data from the appropriate incremental checkpoints. This approach minimizes data loading time for checkout and transparently restores the state in the same kernel process without interruption and with sub-second latency.
@@
Difference from Existing Work
Our work enables high-efficiency time-traveling for computational notebooks through significantly different techniques compared to existing work. While OS-level tools (CRIU, 2023; Ansel et al., 2009) can perform incremental checkpointing, they fail to exploit the fine-grained deltas in data science workflows, cannot perform incremental restore, and fail on remote objects (e.g., Ray (Moritz et al., 2018), on-device data (PyTorch, 2024)). Existing application-level tools (Foundation, 2023g, f; Li et al., [n.d.]) are built for single-time C/R and lack both incremental storage for subsequent checkpoints and incremental restoration featured in this work. Works for variable versioning (Koop and Patel, 2017) and lineage capturing (Brown et al., 2023; Wang et al., 2022) serve significantly different purposes (e.g., visualizing state evolution (Wang et al., 2022)) and cannot be used to travel to a previous/different state, thus are largely orthogonal. Our work shares similarities with time-traveling and versioning in DBMS (Oracle, 2024a; Group, 1996b; Soroush and Balazinska, 2013; Schule et al., 2019); however, our techniques handle complex access patterns and inter-variable dependencies unique to notebook states to enable delta computation for arbitrary objects. We summarize differences in LABEL:tbl:existing_work.
@@ 2. Motivation
This section describes use cases for time-traveling in notebooks (§ 2.1), our intuition for efficient time-traveling (§ 2.2), and how we choose a granularity for incremental checkpointing/checkout (§ 2.3).
@@ 2.1. Why is Time Traveling Useful?
Data cleaning and visualization operations are oftentimes irreversible (e.g., df = df.drop_col(’a’)) and/or partially performed with unobservable side-effects (e.g., a mid-cell execution error), and the user may want to return to the previous state if the execution was erroneous or undesirable (Groups, 2024; Brachmann and Spoth, 2020). To enable interruption-free time-traveling, we can checkpoint the state delta to storage after each operation such that the session state prior to performing the operation can be returned to via loading the appropriate deltas. We empirically study this use case in § 7.5.1.
@@
Reactive execution allows users to quickly investigate alternative paths in data science workflows: when a previously executed cell is re-executed (e.g., train model with different hyperparameters), all its dependent cells (e.g., plotting cells) are too (Koop and Patel, 2017; Shankar et al., 2022). If we can efficiently persist all different variations of objects in different execution paths (i.e., as incremental deltas w.r.t. the shared state), users can efficiently evaluate each variation of values against each other: to switch paths, only the (small portion of) data differing between paths need to be updated via loading the appropriate deltas. We empirically study this use case in § 7.5.2.
@@ 2.2. Enabling Time Traveling
We discuss pros and cons of different checkpointing and checkout approaches for enabling time-traveling to a previous state.
@@
Tools such as CRIU (CRIU, 2023) and DMTCP (Ansel et al., 2009) can be used to create memory snapshots of notebook processes, which contain all data in the session state. (Incremental Checkpointing) Subsequent snapshots can be made incrementally w.r.t. previous snapshots such that only dirty memory pages are stored. However, memory page granularity is coarse and incurs high checkpoint storage overhead (§ 7.3), and OS-level checkpointing is limited to single processes, failing on notebooks utilizing multiple/remote processors (e.g., Spark (Zaharia et al., 2010) and Ray (Moritz et al., 2018) pipelines) or moving data out of CPU (e.g., onto the GPU). (Complete Checkout) Memory snapshots must be entirely loaded to restore the notebook state and additionally require the existing notebook process to be killed (otherwise, a PID conflict occurs) before restoration. This process is both not seamless and incurs high data loading costs (§ 7.5).
Application-level Session Dump
Application-level tools such as Dill’s dumpsession (Foundation, 2023g) and ElasticNotebook (Li et al., [n.d.]) can be used to create checkpoint files by serializing data (e.g., into bytestrings) in the session state. (Complete Checkpointing) Notably, no existing application-level tool supports incremental checkpointing: each checkpoint must independently contain sufficient data for restoring the session state. (Complete Checkout) Checkpoint files need to be entirely loaded (via data deserialization) for session restoration; despite capable of restoring into an existing notebook process, these tools do not utilize variables that are already present within the kernel which may not need to be updated for faster checkout (§ 7.5).
@@
If we can detect the delta between session states at a finer granularity (compared to OS-level tools) and utilize existing data within the kernel to speedup checkout (compared to Application-level dumps), we can achieve efficient and generalized application-level incremental checkpointing and checkout, as follows: (Incremental and Generalized Checkpointing) We utilize application-level information to track state deltas at a finer-than-memory-page-granularity for storage-efficient incremental checkpointing. For generalizability, we can utilize an object’s reduction (i.e., __reduce__ (Foundation, 2023f)) as storage instructions to handle checkpointing multiprocessing and off-CPU workloads. (Incremental Checkout) If we know the contents of the target session state to restore to (e.g.., in terms of a state snapshot (Bernstein and Goodman, 1983)) and accurately compute its difference from the current state, we can perform incremental checkout by only loading and updating data that differ between the current and target state.
Specifically, we compute state differences at the Co-variable granularity, which we describe and motivate in the next session.
@@ 2.3. Tracking State Delta for Time Traveling
Delta detection between states is required for efficient incremental checkpointing and checkout. We discuss pros and cons of the different methods of tracking the state delta.
Dirty Memory Page Tracking
System-level checkpointing tools (CRIU, 2023; Ansel et al., 2009) track state deltas via dirty memory pages. While fast, the delta granularity is insufficient for tracking Python notebook session states, as (1) modifying any object, regardless of size (e.g., x += 1) will cause the entire page (e.g., 4KB) holding the object to become dirty, and (2) Python data structures (e.g., lists) are often constructed and stored in a fragmented manner, leading to simple operations (e.g., mapping elements in a list in-place) creating multiple dirty pages, leading to high incremental checkpoint costs.
@@
Notebook systems such as IPyFlow (Macke et al., 2020), Dataflow (Koop and Patel, 2017), and ElasticNotebook (Li et al., [n.d.]) track session state evolution at the variable-level via provenance-based code analysis (Foundation, 2023a). While capable of producing fine-grained deltas, the tracker has to be either conservative on identifying changed variables (e.g., w.r.t. control flows and external function calls (Li et al., [n.d.])) causing many false positives and large deltas, or perform extensive live instrumentation for resolution (Macke et al., 2020), which can result in high overhead (§ 7.6).
@@
In order to avoid low granularity of memory-page tracking and potential efficiency issues of provenance tracking, we propose to use live object comparison to track updates to Co-variables—connected components of objects (w.r.t. pointer references). Our intuition is that (live) tracking and checkpointing/checking out individual objects are expensive and risky (i.e., may break shared references (Li et al., [n.d.])), respectively. However, at Co-variable granularity, we can achieve low-overhead tracking by reasoning in a principled way based on access patterns which Co-variables were updated by each cell execution, and correctly store/load Co-variables during incremental checkpointing/checkout as if they are independent data tables (§ 4). We depict this idea in LABEL:fig:background_variable_blob: {ls,obj} is a Co-variable (red) as the objects reachable from these variables overlap, i.e., &ls[1] = &obj.foo. {df} is another Co-variable (blue), and there is no way to reach objects under df from objects under ls or obj via references. Notably, Co-variables are the minimum granularity which data in the session state can be stored and loaded without risking breaking shared references (e.g., unlike variable-level KV-stores (shove, 2024; 0xnurl, 2024)). We formally describe Co-variables and how we correctly and efficiently capture state delta at this granularity in § 4.
Motivating Example
Suppose a data analyst is performing text mining (LABEL:fig:background_example). They load the corpus (Cell 1), define category lists (Cell 2), and sort texts by sentiment into the lists (Cell 3). The analyst incrementally checkpoints the state after each cell execution.
(1) Incremental Checkpointing: The analyst wishes to map the lists to clean the contained text and tests with sad_ls (Cell 4, blue). Due to its interleaved construction (with other lists), the list sad_ls is fragmented; incrementally checkpointing the state at memory page granularity for Cell 4 (w.r.t. Cell 3) results in all pages overlapping with sad_ls being copied. However, a Co-variable granularity incremental checkpoint stores only (the bytestring of) sad_ls.
(2) Incremental Checkout: The analyst decided to undo the mapping function in Cell 4 as the results were unsatisfactory. Returning to the state of Cell 3 by (completely) loading a memory snapshot is slow as it requires reloading the corpus. However, by identifying that the states between Cells 3 and 4 differ only by sad_ls, we can only load the value of sad_ls from Cell 3 (red) to replace value of sad_ls from Cell 4 (blue) to perform incremental checkout while keeping the rest of the session state untouched.

### v2

@@ 1. Introduction
Oftentimes, during a workflow, users would like to revert changes made to the session state (i.e., time-travel), such as to undo a modification (e.g., restore a dropped column of a dataframe (StackOverflow, 2024)), restoring an overwritten variable (Groups, 2024), or perform reverse debugging (Brachmann and Spoth, 2020). Unfortunately, unlike program debuggers (e.g., gdb) (Phang et al., 2013; Barr and Marron, 2014; GDB, 2024), relational databases (e.g., PITR in PostgreSQL and MySQL (Group, 1996a; Oracle, 2024b)) or interactive data systems (Dunne et al., 2012; Crotty et al., 2015; Kraska, 2021) which support time-traveling to past program states, existing notebook systems do not natively keep track of past session states: cell executions cannot be undone, e.g., the user cannot ’un-drop’ a dataframe column. If the user executes a cell that alters the session state, a common approach to restore the previous state would be to restart the kernel and then (painstakingly) re-run past cells in the correct order. While code versions can be saved using tools such as Git (Git, 2024b) or native commands (e.g., Jupyter’s %checkpoint (Team, 2023b) [^1] Despite its name, %checkpoint only stores cell code and not objects in the state.) to simplify identifying cells to rerun for restoration, cell reruns can still be time-consuming (e.g., re-training an ML model) and/or result in incorrect restoration (e.g., random train-test splits). Another approach is for the user to periodically checkpoint the session state (e.g., memory dump (CRIU, 2023; Ansel et al., 2009) or session state serialization (Foundation, 2023h; Weinman et al., 2021)) to storage or a managed database (e.g., KV-store (Team, 2023c)). Then, users can load an appropriate checkpoint file to restore the session state. However, performing session checkpointing and restoration using these tools is limiting: checkpointing can incur prohibitive costs (§ 7.3, § 7.4) and may fail on certain workloads (e.g., GPU (cri, [n.d.])), and restoration can either (1) only be (inefficiently) performed from scratch, requiring completely loading a checkpoint file (Foundation, 2023h) and/or killing the current kernel (CRIU, 2023), or (2) may be incorrect, breaking inter-variable relations (Team, 2023c).
@@
We propose Kishu, a notebook system that enables time-traveling between session states: as the user executes cells, Kishu tracks the session state evolution while writing per-cell incremental checkpoints containing differing data between successive states (i.e., the state delta) for returning to any past state via an incremental checkout later. Kishu pursues three challenging goals—Delta-Efficient Checkpoint: Kishu aims to minimize incremental checkpointing overhead by exploiting the small per-cell deltas typical of data science workflows (§ 2.2), but also avoid high detection overhead in the face of complex access patterns and inter-variable dependencies. Correct & Non-intrusive Checkout: Kishu aims to restore past states in the same session non-intrusively by leveraging existing objects in the kernel (that don’t need updating) to minimize data loading costs, while still guaranteeing checkout accuracy as if it completely loaded a checkpoint file. Generalizability: Kishu aims to support checkpointing/checkout for almost all notebook libraries (and/or use cases), of which there is a large variety, e.g., notebooks can perform distributed computing (e.g., Spark (Zaharia et al., 2010)) or move data off-CPU (e.g., GPUs (Foundation, 2024a)). If Kishu can achieve these goals, Kishu will allow users to undo almost any executed cell that undesirably modifies the state as if it never occurred by quickly checking out to the pre-execution state at the cost of minimal workflow overhead.
@@
Our core idea for achieving our goals is to capture the state delta with low overhead, but at sufficiently high granularity using information exclusive to the application level, as follows:
@@
Second, for correct incremental checkout, Kishu identifies the difference between the current and target states at the aforementioned Co-variable granularity according to the checkpoint graph. Then, it replaces (only) Co-variables that need updating in the state by loading data from appropriate incremental checkpoints, minimizing data load time for checkout and transparently restoring the state in the same kernel process without interruption, at sub-second latency.
@@
Difference from Existing Work (LABEL:tbl:existing_work)
Our work enables efficient time-traveling for computational notebooks through significantly different techniques vs. existing work. OS-level tools (CRIU, 2023; Ansel et al., 2009) can incrementally checkpoint notebook states, but fail to exploit the fine-grained deltas of data science workflows (§ 2.2), cannot incrementally restore, and fail on remote objects (e.g. Ray (Moritz et al., 2018), on-device data (PyTorch, 2024)). Existing application-level tools for saving state (Foundation, 2023g, f; Li et al., 2023a; Weinman et al., 2021) lack both the incremental checkpointing and restoration capabilities of our work. Works for versioning notebook cell code (Brachmann et al., 2019; Guo and Seltzer, 2012; Team, 2023b) help identify cells to rerun for restoration but do not directly enable it. Incremental checkpointing/PITR in DBMS (Antonopoulos et al., 2019; Oracle, 2024a)/HPC (Lin et al., 2020; Keller and Bautista-Gomez, 2019) and time-traveling DBMS (Soroush and Balazinska, 2013; Schule et al., 2019) focus on robust and fast logging of table/KV-store updates, while our work focuses on delta computation for complex, interdependent objects unique to notebooks, i.e., what to log (§ 2.2). Orthogonal works include notebook systems for speeding up data exploration using non-time-travel methods (e.g., code recommendation (Lee et al., 2021; Li et al., 2023b)).
@@ 2. Motivation
This section describes use cases for time-traveling notebooks (§ 2.1), characteristics of notebook workloads (§ 2.2), and, accordingly, our intuition for time-traveling (§ 2.3) and capturing state delta (§ 2.4).
@@ 2.1. Why is Time Traveling Useful?
Data operations can be irreversible (e.g., df=df.drop_col(’a’)), and users may want to return to the previous state if the operation outcome is undesirable (Groups, 2024; Brachmann and Spoth, 2020). To enable interruption-free time-traveling, we can checkpoint the state delta to storage after each operation such that the session state prior to performing the operation can be returned to via loading the appropriate deltas. We empirically study this use case in § 7.5.1.
@@
Data scientists often manually create branching, out-of-order cells, where each intended execution path consists of only a subset of notebook cells (e.g., cleaning steps for different models) (Rule et al., 2018; Kery et al., 2018). If we can efficiently persist all variations of objects in different paths as incremental deltas w.r.t. a shared state, users can efficiently switch paths for comparisons: only the (small amount of) data differing between paths need to be updated via loading deltas. We empirically study this use case in § 7.5.2.
2.2. Characteristics of Notebook Workloads
Data scientists commonly use computational notebooks for exploratory work (Kery et al., 2018), which exhibit these key characteristics:
Think Time
Notebook workflows often follow a loop of users sequentially [^2] Existing notebook systems, e.g., Jupyter, do not support concurrent cell executions. writing and running cell code, then observing output (Rule et al., 2018). The notebook kernel can be inactive mid-loop—$\sim$10 seconds of think time (Eichmann et al., 2020) when users decide what cells to write and run next. Notebook systems use think time for computations (e.g., speculative materialization (Xin et al., 2021)). Kishu can also leverage think time for incremental checkpointing.
Complex Access Patterns
Python functions in notebooks can exhibit complex access patterns (e.g., accessing non-parameter variables), making it difficult and/or costly to analyze changes. Hence, some works use a rule-based approach to hard-code effects of common functions (e.g., print not modifying the namespace) (Macke et al., 2020); however, as data scientists also commonly implement and import custom functions into notebooks (Rule et al., 2018), an efficient approach that handles arbitrary functions is highly desired.
Incremental Operations
Users often run cells incrementally, each consisting of only a few lines of code accessing a small number of variables, for fast iteration (Kery et al., 2018). Accordingly, we observe in on of our test notebooks, Sklearn (LABEL:fig:background_workload_characteristics, top), that 40/44 cells access ¡10% of the state data. For these workloads, Kishu’s incremental checkpointing can offer larger benefits.
Data Modifications
Notebook providers often limit session memory consumption (Google, 2024; Kaggle, 2024). Hence, in addition to creating new data, users also modify/delete existing data to conserve memory (e.g., after they have been used to generate relevant cell outputs). This is seen in Sklearn (LABEL:fig:background_workload_characteristics, bottom), where updated data is, on average, split 45%:55% between creations and modifications. Kishu must efficiently undo these in-place updates during checkout.
@@ 2.3. Enabling Time Traveling
We discuss the pros and cons of incremental checkpointing and checkout approaches for enabling time-traveling to a previous state.
@@
Tools such as CRIU (CRIU, 2023) can snapshot notebook process memory to save session state data. (Incremental Checkpointing) Subsequent snapshots can be made incrementally w.r.t. previous snapshots storing only dirty memory pages. However, memory-page granularity is too coarse for incrementally checkpointing notebook workloads as Python data structures (e.g., lists) can be fragmented, leading to simple operations (e.g., in-place sorting a list) causing multiple dirty pages and high checkpoint costs (§ 7.3). (Complete Checkout) Memory snapshots must be entirely loaded to restore a state and additionally require killing the existing notebook process (to avoid PID conflicts) before restoration, which is not seamless and incurs high data loading costs (§ 7.5). Furthermore, OS-level snapshotting is limited to single processes, hence fail on notebooks utilizing multiple/remote processors (§ 7.3).
@@
We can utilize (1) application-level information to detect state deltas at a finer granularity compared to memory snapshots and (2) existing data within the kernel for incremental checkout to address the drawbacks of OS-level snapshotting: (Incremental and Generalized Checkpointing) We track state deltas at a more logical Co-variable granularity (described and motivated in the next session) by tracking references in the notebook for incremental checkpointing. For generalizability, we utilize an object’s reduction (i.e., __reduce__ (Foundation, 2023f)) as storage instructions to checkpoint multiprocessing and off-CPU workloads (§ 6.1). (Incremental Checkout) As we have direct access to the notebook kernel at the application level, we can incrementally checkout by only updating data in the kernel that differ between the current and target states after computing said difference (e.g.., via state snapshots (Bernstein and Goodman, 1983)) (§ 5.2).
@@ 2.4. Tracking State Delta for Time Traveling
We discuss pros and cons of methods of tracking the state delta at the application level for incremental checkpointing and checkout.
@@
Some existing notebook systems (Macke et al., 2020; Koop and Patel, 2017; Li et al., 2023a) track session state evolution via provenance-based code analysis (e.g., via ASTs (Foundation, 2023a)) at a variable-level granularity. Pure static code analysis requirew conservativeness on identifying changed data w.r.t. control flows (e.g., if (x<1):) and external function calls (§ 2.2) causing false positives (e.g., by assuming an untaken branch as taken) and large deltas; hence, these systems augment code analysis with varying degrees of live instrumentation at cell runtime (e.g., resolving the value of x when evaluating the aforementioned if statement) (Macke et al., 2020), which can potentially result in high overhead (e.g., repeated resolutions in looping control flows, § 7.6).
@@
In order to avoid potential efficiency issues caused by the runtime resolution of provenance tracking, we propose to perform live object comparison (i.e., comparing data before and after execution) only after each cell execution to track per-cell updates. Our intuition is that while doing so at variable-level granularity (i.e., like existing provenance trackers) by comparing all objects in the state can be expensive (§ 7.6), it is also unnecessary as storing/loading individual variables (e.g., with variable-level KV-stores (shove, 2024; 0xnurl, 2024)) during checkpointing/checkout risks breaking shared references (Li et al., 2023a). We hence track updates at a coarser Co-variable level—connected components of objects (w.r.t. pointer references): we can reason based on access patterns which Co-variables were possibly/definitely not updated by each cell to limit object comparisons, and correctly store/load Co-variables as if they are independent data tables. LABEL:fig:background_variable_blob depicts our idea: {ser,obj} is a Co-variable (red) as the reachable objects of the variables overlap (&ser[1] = &obj.foo). {df} is another Co-variable (blue), and users cannot reach objects under df from objects under ser or obj via references. Notably, Co-variables are the minimum granularity which data in the state can be stored/loaded without risking breaking shared references. § 4 formally describes Co-variables and how we correctly and efficiently capture state delta at this granularity.
Motivating Example (LABEL:fig:background_example)
Suppose a data analyst is performing text mining. They load the corpus (Cell 1), define category lists (Cell 2), and sort texts by sentiment into the lists (Cell 3). The analyst incrementally checkpoints the state after each cell execution. (1) Incremental Checkpointing: The analyst tests a mapping to clean the contained text in the lists on sad_ls (Cell 4, blue). Due to its interleaved construction (with other lists), the list sad_ls is fragmented; incrementally checkpointing the state at memory page granularity for Cell 4 (w.r.t. Cell 3) results in all pages overlapping with sad_ls being copied. However, a Co-variable granularity incremental checkpoint stores only (the bytestring of) sad_ls.
(2) Incremental Checkout: The analyst undoes the mapping in Cell 4 due to poor results. Returning to Cell 3’s state by (completely) loading a memory snapshot is slow as it requires reloading the corpus. However, by identifying that the states of Cells 3 and 4 differ only by sad_ls, we can only load the value of sad_ls from Cell 3 (red) to replace value of sad_ls from Cell 4 (blue) to perform incremental checkout without touching the rest of the session state.